OpenStudy (anonymous):

the perimeter of a standard sized rectangular rug is 24 ft. The length is 2 ft longer than the width. Find the dimensions. what is the width? what is the length?

OpenStudy (marigirl):

we can say the width is 'x', then the length is x+2 then we can write the perimeter as x+x+x+2+x+2=24 4x+4=24 can u go from here?

OpenStudy (anonymous):

no I don't get it

OpenStudy (marigirl):

would u like a diagram?

OpenStudy (marigirl):

i can draw it

OpenStudy (anonymous):

surre

OpenStudy (anonymous):

no it wont open it

OpenStudy (marigirl):

okay, lets say the width of this rectangle is x .. an unknown then the length is 2ft longer than width .. so we can write it as width+2 which is x+2 are u with me so far

OpenStudy (anonymous):

yes

OpenStudy (marigirl):

cool, now the perimeter is the measurement around the outside.. we will have two sides of width and two sides of length right?

OpenStudy (anonymous):

yes

OpenStudy (marigirl):

so its like writing x + x + x+2 + x+2=24 (Both widths) (both lengths) (total of 24ft given in question)

OpenStudy (anonymous):

okay so how would I do it

OpenStudy (marigirl):

then we solve for x.. there are 4 'x' on the left side and then 2+2 4x+2=24 4x=20 x=5.. which means the width is 5

OpenStudy (marigirl):

dont forget the length is 2ft longer than width

OpenStudy (anonymous):

so width is 5 and length is 7

OpenStudy (marigirl):

Yes :)